The velocity ( v meters per second ) of a body is known to be proportional to the sqaure root of it's kinetic energy ( e ejoules). When the velocity of a body is 120m/s it's kinetic energy is 1600J.
Write down the relationship between v , e using k as the constant of varation.
Calculate the value of k
v = 21 , calculate the kenetic energy of the body in joules.

Can someone explain me please the right method in solving each quesion with the steps.

Answer by edjones(8007)

v=k*sqrt(e)
120=k*sqrt(1600)
120=40k
k=120/40
k=3
.
21=3*7
e=7^2=49 e joules
.
Ed